OpenStudy (arihangdu):

Verify the identity. (1+tan^2u)(1-sin^2u)=1

OpenStudy (mathmale):

Try this: substitute (cos 2u)^2 for (1-(sin 2u)^2).

OpenStudy (mathmate):

hint: use identities to help you: \(1+tan^2(x)=sec^2(x)\) \(sec(x)=\frac{1}{cos(x)}\) \(sin^2(x)+cos^2(x)=1\)

OpenStudy (welshfella):

these 2 will be useful - the one mention by mathmale and also tan^2 u = sin^2 u / cos^2 u